### v1

# The Inhibitor: ReLU and Addition-Based Attention for Efficient Transformers

Abstract
To enhance the computational efficiency of quantized Transformers, we replace the dot-product and Softmax-based attention with an alternative mechanism involving addition and ReLU activation only. This side-steps the expansion to double precision often required by matrix multiplication and avoids costly Softmax evaluations but maintains much of the core functionality of conventional dot-product attention. It can enable more efficient execution and support larger quantized Transformer models on resource-constrained hardware or alternative arithmetic systems like homomorphic encryption. Training experiments on four common benchmark tasks show test set prediction scores comparable to those of conventional Transformers with dot-product attention. Our scaling experiments also suggest significant computational savings, both in plaintext and under encryption. In particular, we believe that the ReLU and addition-based attention mechanism introduced in this paper may enable privacy-preserving AI applications operating under homomorphic encryption by avoiding the costly multiplication of encrypted variables.
1 Introduction
Transformer models [1] have achieved great success in various machine-learning tasks and are the foundation for modern large-scale pre-trained Language Models such as BERT [2], RoBERTa [3], XLNet [4], Transformer-XL [5], and the GPT family [6, 7]. These models have enabled leaps in performance on many NLP tasks, as for example experienced by the millions that use the ChatGPT application [8]. Similarly, the Vision-Transformer (ViT) [9], and similar models outperform CNNs on many computer vision tasks such as image classification [10] and object detection [11].
The impressive gains in performance, however, come at a price. State-of-the-art Transformers are extremely large, having parameters in the hundreds of billions. The requirements for memory and computational power have become prohibitive, not only for deployment on resource-constrained embedded systems, but sometimes even in data centers, and their considerable energy usage is widely discussed in society. Neural network quantization [12, 13, 14] is a method to address this problem, as it reduces memory usage by employing lower-bit precision for weight and activation tensors. This not only decreases inference time but also increases energy efficiency through the adoption of low-bit fixed-point arithmetic instead of floating-point operations [15]. Other approaches to improve Transformer’s scalability and environmental sustainability include Pruning [16, 17], Knowledge Distillation [18, 19] and Neural Architecture Search [20]. In this work, we propose changes to the architecture’s fundamental building blocks to take steps towards the same end.
Transformers use an attention mechanism to identify connections between different elements in a sequence, which conventionally takes the dot-product between the query and key matrices before passing the results through a Softmax function. Many alternative formulations have been proposed, such as the Fastformer [21] based on additive attention, or the ReLUformer [22] that uses ReLU activation in place of Softmax. These architectures share elements with what is proposed in this work, however, they still apply attention by a matrix multiplication with the value matrix. Extending our earlier work with ReLU and addition-based gated RNNs [23], we here propose the Inhibitor attention mechanism, which is designed to not rely on variable-to-variable multiplication and Softmax activation. These operations are more costly than constant-to-variable multiplications (also known as literal multiplication) and ReLU on many computing hardware. Even more so, they are particularly challenging operations under homomorphic encryption, an arithmetic system that permits calculation with encrypted variables without access to the secret key [24].
2 Method
The Transformer architecture consists of repeated blocks of self-attention and feed-forward networks (FFNs). In the simple decoder application, each block takes as input a sequence embedding, $X\in\mathbb{R}^{n\times d}$, where $n$ is the length and $d$ is the dimension. The embedding is transformed into query, key and value, by multiplication with weight matrices according to $Q=XW_{Q}$, $K=XW_{K}$, and $V=XW_{V}$. Attention scores are calculated by passing the scaled dot-product of query and key through the Softmax function,
$S=\mathrm{Softmax}\left(QK^{T}/\sqrt{d}\right)$ (1)
which is then used to form a weighted sum of values as output, $H=S\,V$, exploiting that each row of $S$ is normalized to sum to one. The FFN has two fully connected layers with weight matrix multiplication and ReLU activation,
$H=(XW_{1}^{T}+b_{1})^{+}W_{2}+b_{2}$ (2)
where $W_{1},W_{2}$ are weight matrices and $b_{1},b_{2}$ are bias vectors. We use the short-hand notation $x^{+}=\max(0,x)$ for the ReLU function. Each Transformer block typically also has one or more layer normalizations [25] that stabilize the gradient flow and contribute to regularizing the training.
The Inhibitor.
We propose to alternatively calculate attention scores as
$Z_{ij}=\sum_{k}\frac{1}{\gamma}\left|Q_{ik}-K_{jk}\right|$ (3)
where mixing between query and key is based on their absolute difference instead of dot-product. We have thus replaced the cosine (dot-prod) distance of conventional attention with the Manhattan distance. In place of Softmax multiplication, we subtract attention scores from values inside a ReLu function,
$H_{ik}=\sum_{j}\left(V_{jk}-Z_{ij}\right)^{+}$ (4)
such that entries of $V$ for which $Z$ are large are zeroed out and do not contribute to the sum. We call this mechanism inhibition as it is reminiscent of subtractive inhibition in biological neurons. FFN and normalization are left unchanged.
Shifted inhibition score. The Manhattan score of equation 3 can only take the value zero for rows where $Q$ and $V$ are identical. To make it easier to obtain a zero inhibition score, which allows entries of the value matrix, $V$, to pass unmodified through equation 4 we test using a slightly modified inhibition score, $Z^{\prime}=(Z-\alpha)^{+}$, which applies a constant shift $\alpha\geq 0$ to the Manhattan score.
3 Results
Benchmark comparisons.
We carried out numerical experiments that trained Transformer models based on the new Inhibitor and the conventional dot-product attention on four standard tasks. The aim was not to achieve SotA results but rather to examine if the Inhibitor mechanism would perform comparably on a set of familiar tasks, which is why we used simple set-ups without hyperparameter tuning. From Table 1, which reports the results, we note that for each task, the two alternative attention mechanisms score very similarly. Indeed, none of the differences are significant at 95% confidence (over at least 20 repetitions for each experiment). For experiments with the Inhibitor, we used a shifted score with $\alpha=0.5$ and scaling factor $\gamma=\sqrt{d}$.
[…]
all_results.txt
Table 1: The Inhibitor shows comparable performance to conventional attention for Transformers trained on four standard tasks (for mse, acc, acc, and edit distance, respectively).
The adding problem was introduced by [26] to test long-term memory for RNNs. It takes two inputs: a sequence of random numbers and a two-hot sequence (which, in our experiments, each is 100 long). The ground truth is simply the dot-product of the two inputs as vectors, which, rather obviously, is not a challenging task for a conventional (dot-prod-based) Transformer. It is encouraging that the addition-based Inhibitor performs just as well on this task (which is hard for, e.g., simple RNNs to solve).
We also train one-layer Transformers for the MNIST handwritten digit recognition task [27] and IMDB movie-review sentiment analysis task [28], which are simple go-to benchmark task for image classification and text analysis, respectively. Although far from SotA, we note that our simple models achieve decent accuracy for both attention mechanisms, where the differences in performance are not significant.
The final task, labeled IAMW in Table 1, uses the IAM Handwriting Database [29], which is a collection of handwritten words written by more than 700 writers. The model first applies two convolutional layers followed by a Transformer and uses Connectionist Temporal Classification (CTC) loss [30] as an endpoint layer to predict the text. Edit distance [31] was used as the evaluation metric. Again, the differences are small.
Scaling experiments.
Next, we wanted to examine and compare the scaling properties for the proposed mechanism under two identified scenarios: i) quantization with integer arithmetics and ii) homomorphic encryption. Therefore, the two alternative attention mechanisms were implemented directly in low-level code rather than high-level ML libraries, where built-in optimizations for conventional models and design choices would bias a comparison.
For the plaintext experiment, we used integer 16-bit arithmetics implemented in the Rust programming language, which gives detailed low-level control over circuits and supports advanced time benchmarking through the Criterion package. The encrypted implementation uses the TFHE scheme [32] for Fully Homomorphic Encryption on the Torus as supported in the Concrete library for Python [33]. TFHE permits the evaluation of arbitrary single-variate activation functions under homomorphic encryption through the Programmable Bootstrap [34]. We here used considerably smaller networks where the embedding dimension was limited to size 2. The encrypted circuits were implemented for integers with up to 8-bit precision.
[…]
timing_short_plain.txt
Table 2: Estimated plaintext execution time on CPU for four different sequence lengths (with fixed size single head).
[…]
timing_short_tfhe.txt
Table 3: Estimated encrypted execution time on CPU for four different sequence lengths (with fixed size single head).
The results indicate that the proposed Inhibitor mechanism can have a significant advantage, with i) 30%–50% saving for the plaintext implementation on CPU as per Table 2, and ii) a factor 3–6 under encryption with TFHE as per table 3. Timing estimates are averaged over 20 repeated experiments and are significant at the 95%
Conclusions.
The proposed Inhibitor mechanism allows straightforward quantization as equations 3 and 4 naturally support integer implementation. Our scaling experiments indicate that it has considerable potential for computational savings both for Transformers processing plaintext data as well as ciphertext.
While experiment results are promising on simple training tasks, for future work, it is necessary to examine performance under more challenging settings, for example, by pre-training a much larger Transformer model and testing on modern NLP and image recognition benchmark tasks.
Appendix
3.1 Signed Inhibitor
The inhibition mechanism described in equation 4 can only produce a non-negative attended value matrix $H$, while the conventional dot-product attention can also select from $V$ that can have both positive and negative entries. We can rewrite inhibition to handle signed values simply as
$H_{ik}=\sum_{j}\left(V_{jk}^{+}-Z_{ij}\right)^{+}+\sum_{j}\left(V_{jk}^{-}+Z_{ij}\right)^{-},$ (5)
such that the ReLU terms within the sum will attenuate or extinguish $V_{jk}$ for large $Z_{ij}$, while $V_{jk}$ will pass through unaltered for $Z_{ij}=0$.
3.2 Implemention
A naive implementation of equation 3 and 4 relies on expanding the dimension of the broadcasted differences before passing the result through an $\mathrm{abs}$ or $\mathrm{ReLU}$ function and then summing over an inner dimension. This uses large amounts of memory and should be avoided.
The ML software that we used for this study supports fused operations for certain operations that avoid memory bloat and make the execution of the inhibition much more efficient on CPU and GPU hardware. Specifically, it supports calculating the pairwise distance between two tensors [^1] The cdist function in the PyTorch library for deep learning., which directly implements calculating the Manhattan distance of equation 3.
For the inhibition step of equation 4, we use a simple trick based on
$x^{+}=\frac{x+|x|}{2}$ (6)
and write
$H_{ik}$ $=\sum_{j}\left(V_{jk}-Z_{ij}\right)^{+}$
$=\frac{1}{2}\sum_{j}V_{jk}-\frac{1}{2}\sum_{j}Z_{ij}+\frac{1}{2}\sum_{j}|V_{jk}-Z_{ij}|$ (7)
where the first two terms are simple sums that do not expand memory usage, and the last term can benefit from efficient fused pairwise Manhattan distance implementations.
Similarly, we rewrite for the signed inhibition of equation 5 to have
$H_{ik}$ $=\sum_{j}\left(V_{jk}^{+}-Z_{ij}\right)^{+}+\sum_{j}\left(V_{jk}^{-}+Z_{ij}\right)^{-}$
$=\frac{1}{2}\sum_{j}V_{jk}+\frac{1}{2}\sum_{j}|V_{jk}^{+}-Z_{ij}|-\frac{1}{2}\sum_{j}|V_{jk}^{-}+Z_{ij}|,$ (8)
where we have, in addition to equation 6, used the relation
$x^{-}=\frac{x-|x|}{2}$ (9)
for the negative ReLU function to leverage a fused implementation.